AGw Commissioning with IRTC 08-09 November 2008 MST

Observers: A. Rakich, D. Miller (at beginning) in Tucson, Nancy in Germany
Operator: A. Ceranski

Summary

Details (Times reported in MST below)

5:15 Aaron turned on all systems, opened the chamber, and sent a preset with no problems, failures or errors!

5:50 Sent preset to M5_0835. No star. PCS had been restarted so default IE=45 and CA=-69. Set IE to last night's value of -20 and the star was on the IRTC and within 10" of the center of rotation. Set IE=-30, CA=-67 .

5:59 pointto BS9188 TelMode="Active" gstar=0 Success!

6:00-6:37 Collimate WFSC 1-23 WFE<350nm in 6 iterations (first three were 5 sec WFSC exposures because we had shortened for Titan last night). WFE<200nm for last 10 iterations.

6:37 pointto BS9183 TelMode="Active" gstar=0 Success! On-axis star is very dim ~200 counts. Think we acquired a nearby dim star.

6:53 pointto M5_1526 TelMode="TRACK" Star 15'+ low and to the right. Set IE=-45 CA=-62 to center. Want star where there will be minimal elevation change so we can measure Z11-Z4 scale and Z22-Z4

6:56 pointto BS9183 TelMode="ACTIVE" gstar=0 Success! Bright star on-axis. 3500 counts on WFSC

6:56-7:13 collimate gcs 36-51 Seeing now ~0.5"

7:14 IDL agw_collimate (do not send) wfsc 01, wfsc 02 (send collimate)

7:16 put Z11 +1000 on primary, CLUT, wfsc 03, CLUT, 04, CLUT, 05 (Z4 = -431, ?, 337)

7:19 Z(act_opt) = 0.252, put Z11 = -1000, Z(act_opt) = 0.125, CLUT, 06, CLUT, 07

7:21 put Z11 +1000 on primary, Z(global) + 0.02 on primary, CLUT 08, clut, 09, CLUT, 10 (see ~-133 Z4)

7:27 put Z(global) +0.029 on primary, CLUT, 11, CLUT, 12 (see ~60 Z4)

7:29 remove Z11 and Z4 Collimate with gcs 52-56

7:35 Send focus 0 to offload secondary hexapod

7:36 gcs 57 ~100nm WFE 58 190nm WFE

7:38 put Z11 +1000 on primary, Z(global) + 0.029, wfsc 13, CLUT, 14, CLUT, 15 (see 368, 452 Z4)

7:40 Z(global) + 0.015, wfsc 13 (see 330 Z4) will recollimate

7:44 remove z11 and z4

7:44 collimate gcs 59-64

7:50 put Z11 +1000 on primary, Z(global) + 0.015, wfsc 16, CLUT, 17, CLUT, 18 (see -1.7, -113, )

7:52 Z(act_opt) = 0.256 remove Z11 Z(act_opt) = 0.110, remove Z4

7:52 GCS GUI froze. Had to force quit.

7:58 Started new GCS GUI. Not responding

8:00 pointto BS183 TelMode="Active" gstar=0 Success!

8:01 ctrl-C GCS Gui. Start new GUI. Still not repsponding. Will use Guider2 python script instead

8:10 Andrew Take Z22 data.

8:25 last collimation file wfsc 89, send +100 nm of Z22 with FCin set to zero, looking for -2290 nm Z4. Not there, so PSF is not set up to work for Z22 as it does for Z11 yet.

8:33 noticed that with 100 nm Z22 sent the scale factor was only around 2, much less than when we were trying to send 500 nm Z22. Checking now for linearity.

8:42 restarted GCS and now the GUI is working again. Going to new star BS9101

8:50 last collimation in GCS wfsc 98 in /tmp, now swending +100 Z22 and measuring 3 sets in /AGW_Data/081109w

8:53 screwed up, startting again, collimating in /AGW(using IDL) wfsc 26-27 then focus 0 in IRAF to offload M2 corrections to M1, measure next file wfsc 28, get a large focus (500 nm ) with the offload. Needs to be investigated.

8:56 collimating wfsc 28-29. Send +100 Z22, wfsc 30-32, Z22 before adjustment 62 nm, after adjustment Z22 -> 90nm, 90 nm, 84 nm, so scale factor of ~3.3 confirmed from last weeks measurements.

9:04 investigating focus piggybacking with M2 active optics offload; clear all active optics from both mirrors and collimate in IDL wfsc 35- 43

9:15 collimated, active Z on M2 is -0.227, step focus measure wfsc screwed up again, should have noted active Z on both M1 and M2 first. start again collimating with alll active optics removed wfsc 44-51

9:23 active z M1 0.072, M2 -0.217 step focus, M1-> .218 M2 ofload worked, now on the M1 gui showing up as +0.218 mm. wfsc 52-54 show good correctionm, must have been a "blip" we saw before (8:53 tonight).

9:29 going to test GCS/IDL parity and convergence at edge of field now, presetting to BS9010 GS 7 and collimating GCS wfsc 99-108, collimated nicely and pupil in right position, but not so far off-axis (y = 473 mm). Resending preset with PA +45 (to 60 degrees) (x = -97 mm, y = 513 mm) and re-collimating wfsc 109-116 , good collimation (and using new low order scale factors, seems to give slightly better convergence, impression only at this point). Pupil seemed well positioned in both cases.

10:10 sending new preset to BS9107 collimate on GS 0 gcs wfsc 120-121 (corrected from 1000 nm z4-z11 to 90 nm z4-z11 in one step, further addingto positive impression of new coefficients, going to GS 18 Y= 425 mm. GCS wfsc 123-130 Pupil correctly positioned.

10:25 changing AIP_L.cfg focus from 36.6 to 31.6 mm.

10:27 resending preset to GS 18 with PA original + 45 degrees (346 degrees) says it found the star, and a bright star not too far off in acquisition image #13, but not offset ot guiding correctly. manually drove it on, now guiding, collimating wfsc 133-140

10:35 resending preset with PA original + 60 (01) collimating wfsc 141-158, xy (-161, 518).

10:54 resending preset with PA original -60 (241) 10:57 acquisition image 17, clear GS but not guiding (Torsten?). manually fluked it driving onto guide box with IE CA. collimating wfsc 159- 163, looking good.

11:03 sent guide preset on axis and measuring aberration in /AGW_Data/081109, wfsc 58-60 note ~400 nm Z4 offset from GS 18 to GS 0, other aberrations Z5-Z8 very good, perfect actually.

11:04 conclusion, the pupil fit algorithm is now much improved, and GCS is no longer diverging in its corrections at edge of field as a result.

11:07 taking field data on BS9107. collimating on-axis, measuring off-axis, sampling 8 stars in the field. collimating usng IDL, collimating on-axis.

11:14 reset IE -30 CA -62 collimating on axis with GCS wfsc 164-172 (from now on going to send guide presets and collimate with IDL (PS note that way back when collimating in IDL wfsc 63 had an interesting pupil for vibration)

11:28 a cell of bad seeing disrupted proceedings, maybe the first of the front. collimating on-axis with IDL, wfsc 65-67, going to GS 18, measuring 68-70 Doug's angle = -403, y= 426 mm

11:35 back on axis collimating wfsc 68- 72, going to GS 82 and measuring wfsc 73-75 Doug's angle = -493 , y = 485.6

11:42 back on axis collimate wfsc 76-82 going to GS 97 bu tpreset failked rotator fault going back on-axis to collimate... wfsc 83-86 (86 pupil image very distorted but good data)

11:59 taking data on GS 97 wfsc 87-89 dougs angle = -240 , y = 439.8

12:01 collimating on-axis wfsc 90-91, going to GS 39 wfsc 92-94 y = 476.6, DA = -349

12:08 collimating on-axis wfsc 95 going to GS 11 wfsc 97-99, Doug's angle -394, y = 440

12:12 collimating on-axis wfsc 100-110 pause this operation as we are near zenithg, preset to BS 9108 for tracking test.

12:25 pointto BS9108 collimate gcs ->177

12:34 pointto BS9108 Track

12:40 take 15 minutes worth of tracking data on IRTC takeseq 0.5 1800 file in /Repository/IRTC_Data/081109/tracking_00040_cube.fit elevation ~ 64 degrees.

12:50 updated collimation around image # 1170

1:00 slewing back to BS9107 to continue field aberration measurments...

1:05 collimating on-axis wfsc 112-115, go to GS 18 "transfer of initial correction to PCS failed" acquisition image 36 (MICHELE) 1:07.

1:07 back on-axis to collimate... wfsc 116-121, going back to GS 18 wfsc 122-124 Doug's angle -182 y = 426 mm

1:20 back on-axis collimate pointing drift, re-sending preset in track mode and dialing IE CA. IE=-22 CA=-57

1:24 back on-axis to collimate wfsc-125-126 going to GS 70 wfsc 127-129, Doug's angle = -254, y = 457

1:29 back on axis to collimate wfsc 130-133 going to GS 96 wfsc 134-136 Doug's angle = -310, y= 457 mm

1:35 back on axis to collimate wfsc 137- (in the middle of this did a focus 0 to offload M2 corrections and it didnt work at 1:42 tried again and it worked but next wfsc image had 1000 nm Z4 like before tonight see wfsc 144) -168

2:15 going to GS 97 wfsc 169-171 y= 440 mm, Doug's angle = -386, going back on axis to collimate wfsc 172-173

2:22 Going to GS 64 wfsc 174-176 y = 445.7, Doug's angle = -442, going back on axis to collimate wfsc 177-179

2:30 going to GS 19 wfsc 180-182, y = 424, Doug's angle = -502, going back on axis to collimate wfsc 183-186

2:39 going to GS 11 wfsc 187-189, y = 440 , Doug's angle = -532, going back on axis to collimate wfsc 190-195

2:48 going to GS 40 wfsc 196-198, y=441, Doug's angle = -233, going back on axis collimating wfsc 199-200

2:54 go to GS 70, wfsc 201-203, y = 457, Doug's angle = -276, going back on axis to collimate wfsc 204-207

3:01 going to GS 96, wfsc 208-210, y= 457, Doug's angle = -330, going back on axis to collimate wfsc 211

3:06 going to GS 97wfsc 212-214, y= 439. Doug's angle = -394, going back on axis collimate wfsc 215-216

3:12 going to GS 64 wfsc 217-219, y= 446 Doug's angle = , going back to on axis collimate wfsc 220-221

3:20 going to GS 19 wfsc 222-224, y=424 , Doug's angle = -509, going back on axis collimate wfsc 225-228

3:28 going to GS 8 wfsc 229-231, y=432, Doug's angle = -187, enough of this!

Going on to next test: collimating on off-axis stars and measuring on-axis with IRTC.

3:30 pointto M5_0258 ACTIVE gstar=4

3:51 well collimated gcs through 196

3:51 Check exposure and focus for EF images +-0.7, 0.2 (150)

3:55 collimate 197-203

4:00 Start taking EF data. All exptimes are 0.2 with 150 frames, J Band Filter

gstar ef_in ef_ex Probe x,y Doug's angle (gsc#, rms Z4-11
4 42 41 -0.08,143.92 -221 (203,114)
3 43 44 -0.25, 129.25 -135 (204,165) (205,319) (206,289) (207,99)
10 45 46 0.0,167.51 -60 (208,400) (209,225) (210,138)
9 47 48 0.34, 165.63 25 (213,154) (214,155)
6 49 50 -0.2,145.29 -255 (215,602) (216,237) (217,140)
8 51 52 -0.56,155.91 -176 (218,417) (219,125)
12 53 54 0.11,188.54 -36 (220,1035) (221,302) (222,176)
13 55 56 0.09,193.52 -78 (223,365) (224,277) (225,201)
2 PA=264 57 58 -0.12,121.34 -232 (226, 1243) (227,988) (228,158)
2 PA=204 59 60 -105.16,60 -169 (229,640) (230,125) (231,not sent)
2 PA=324 61 62 -104.96,60 -287 (232,914) (233,391) (233,256) (234,256) (235,233)
2 PA=324 63 64 -104.96,60 -287 Put Z22 +200 on Primary

*Note: at 4:25 guiding found the wrong star (ignore gcs 211 and 212), reset *IE=-32, CA = -63

-- DougMiller - 09 Nov 2008
Topic revision: r8 - 13 Nov 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback